OpenStudy (anonymous):

Solve 4x2 = 28 by finding square roots

jimthompson5910 (jim_thompson5910):

First divide both sides by 4. Then take the square root of both sides (don't forget about the plus/minus) \[\Large 4x^2 = 28\] \[\Large x^2 = \frac{28}{4}\] \[\Large x^2 = 7\] \[\Large x = \pm\sqrt{7}\] \[\Large x = \sqrt{7} \ \text{ or } \ x = -\sqrt{7}\]

OpenStudy (anonymous):

thanks